OpenStudy (anonymous):

Determine the nature of the roots for the following equation. 1/x - 1/x + 1 = 1/5

OpenStudy (experimentx):

1/x - 1/(x + 1) = 1/5 ???

OpenStudy (anonymous):

1 over x - 1 over x + 1 = 1 over 5 (:

OpenStudy (experimentx):

5(x+1 - x) = x(x+1) => 5 = x^2+x => x^2+x-5 = 0 the nature of roots of quadratic equation is determined by 1^2+4*5 > 0 ==> hence there will be 2 roots and real

OpenStudy (anonymous):

Thx (: